# Approximation of multipartite quantum states: revised version with new applications

### v1

@@ 5.3 On the energy-constrained versions of $E_{R,\pi}$
$\mathfrak{S}_{m}(\mathcal{H}_{A_{1}..A_{n}})\doteq\left\{\rho\in\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})\,|\,S(\rho_{A_{1}}),...,S(\rho_{A_{m}})<+\infty\hskip 1.0pt\right\}$
such that
$-c^{-}_{f}C_{m}(\rho)\leq f(\rho)\leq c^{+}_{f}C_{m}(\rho),\quad\quad C_{m}(\rho)=\sum_{s=1}^{m}S(\rho_{A_{s}}),$ (60)
for any state $\rho$ in $\mathfrak{S}_{m}(\mathcal{H}_{A_{1}..A_{n}})$ and
$-d^{-}_{f}h_{2}(p)\leq f(p\rho+(1-p)\sigma)-pf(\rho)-(1-p)f(\sigma)\leq d^{+}_{f}h_{2}(p)$ (61)
for any states $\rho$ and $\sigma$ in $\mathfrak{S}_{m}(\mathcal{H}_{A_{1}..A_{n}})$ and any $p\in(0,1)$, where $h_{2}$ is the binary entropy (defined after (2)), $c^{\pm}_{f}$ and $d^{\pm}_{f}$ are nonnegative numbers such that $c^{-}_{f}+c^{+}_{f}=C$ and $d^{-}_{f}+d^{+}_{f}=D$.
Let $\widehat{L}^{m}_{n}(C,D)$ be the class containing all functions in $L^{m}_{n}(C,D)$ and all the functions of one of the forms
$f(\rho)=\inf_{\lambda}f_{\lambda}(\rho)\quad\textrm{and}\quad f(\rho)=\sup_{\lambda}f_{\lambda}(\rho),$
where $\{f_{\lambda}\}$ is a family of functions in $L^{m}_{n}(C,D)$.
Let $N^{m}_{n,s}(C,D)$, $s=1,2,3$, be the class of all functions $f$ on the set $\mathfrak{S}_{m}(\mathcal{H}_{A_{1}..A_{n}})$ defined by one of the expressions
$f(\rho)\doteq\inf_{\hat{\rho}\in\mathfrak{M}_{s}(\rho)}h(\hat{\rho})\quad\textrm{and}\quad f(\rho)\doteq\sup_{\hat{\rho}\in\mathfrak{M}_{s}(\rho)}h(\hat{\rho})$
via a particular function $h$ in $\widehat{L}^{m}_{n+l}(C,D)$ for some $l>0$, where:
  • $\mathfrak{M}_{1}(\rho)$ is the set of all extensions of $\rho$ to a state of $A_{1}...A_{n+l}$;
  • $\mathfrak{M}_{2}(\rho)$ is the set of all extensions of $\rho$ having the form
    $\hat{\rho}=\sum_{i}p_{i}\rho_{i}\otimes|i\rangle\langle i|,$ (62)
    where $\{\rho_{i}\}$ is a collection of states in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$, $\{p_{i}\}$ is a probability distribution and $\{|i\rangle\}$ is an orthonormal basis in $\mathcal{H}_{A_{n+1}}$ (in this case $l=1$);
  • $\mathfrak{M}_{3}(\rho)$ is the set of all extensions of $\rho$ having the form (62) in which $\{\rho_{i}\}$ is a collection of pure states in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$.
Below is a (incomplete) list of characteristics of a quantum system $A_{1}...A_{n}$, indicating the classes to which they belong with reference to the justification:
  • the von Neumann entropy $S(\rho)$ of a state $\rho$ of $A_{1}$ ($L^{1}_{1}(1,1)$, the concavity of $S$ and inequality (2));
  • the (extended) quantum conditional entropy $S(A_{1}|A_{2})$ defined in (3) ($L^{1}_{2}(2,1)$, [33, Section 4.1]);
  • the quantum mutual information $I(A_{1}\!:...:\!A_{n})$ ($L^{n}_{n}(1,n)$ and $L^{n-1}_{n}(2,n)$, [33, Section 4.2]);
  • the (extended) quantum conditional mutual information $I(A_{1}\!:...:\!A_{n}|A_{n+1})$ ($L^{n-1}_{n+1}(2,n)$ and $L^{n}_{n+1}(2-1/n,n)$, [33, Section 4.2]);
  • the one-way classical correlation in $A_{1}A_{2}$ with measured system $A_{1}$ ($\widehat{L}^{1}_{2}(1,2)$, [33, Section 4.3]);
  • the one-way classical correlation in $A_{1}A_{2}$ with measured system $A_{2}$ ($\widehat{L}^{1}_{2}(1,2)$, [33, Section 4.3]);
  • the quantum discord in $A_{1}A_{2}$ with measured system $A_{1}$ ($\widehat{L}^{1}_{2}(1,2)$, [33, Section 4.3]);
  • the quantum discord in $A_{1}A_{2}$ with measured system $A_{2}$ ($\widehat{L}^{1}_{2}(2,2)$, [33, Section 4.3]);
  • the constrained Holevo capacity of the partial trace over $A_{2}$ in $A_{1}A_{2}$ ($N^{1}_{2,2}(1,2)$, [33, Section 4.5]);
  • the constrained Holevo capacity $\bar{C}(\Phi,\rho)$ of a quantum channel $\Phi:A_{1}\rightarrow B$ at a state $\rho$ of $A_{1}$ ($N^{1}_{1,2}(1,2)$, [33, Section 4.5]); [^21] This follows, due to relation (74) in [34], from the fact that the constrained Holevo capacity of the partial trace over $E$ in a bipartite system $BE$ belongs to the class $N^{1}_{1,2}(1,2)$ in the setting $A_{1}=BE$ [33, Section 4.5].
  • the mutual information $I(\Phi,\rho)$ of a quantum channel $\Phi:A_{1}\rightarrow B$ at state $\rho$ of $A_{1}$ ($L^{1}_{1}(2,2)$, [29, Section 4.3]);
  • the coherent information $I_{c}(\Phi,\rho)$ of a quantum channel $\Phi:A_{1}\rightarrow B$ at state $\rho$ of $A_{1}$ ($L^{1}_{1}(2,2)$, [29, Section 4.3]);
[…]

### v2

@@ 4.2.2 Continuity and convexity properties of $R$ and $R^{\infty}$
$\widetilde{\mathfrak{S}}_{II}^{1}(\mathcal{H}_{AB})=\left\{\hskip 1.0pt\rho\in\mathfrak{S}(\mathcal{H}_{AB})\,|\,\textit{either }\rho_{A}\!\!\in\mathfrak{S}_{\rm\!\textsf{FA}\!}(\mathcal{H}_{A})\textit{ or }\rho_{B}\!\!\in\mathfrak{S}_{\rm\!\textsf{FA}\!}(\mathcal{H}_{B})\hskip 1.0pt\right\}$ (86)
be the subset of $\,\mathfrak{S}(\mathcal{H}_{AB})$ consisting of states $\rho$ such that the FA-property holds for at least one of the marginal states of $\rho$.
A) The functions $R$ and $R^{\infty}$ are finite and lower semicontinuous on the set $\,\widetilde{\mathfrak{S}}_{II}^{1}(\mathcal{H}_{AB})$. Moreover,
$\liminf_{k\to+\infty}R^{*}(\rho_{k})\geq R^{*}(\rho_{0}),\quad R^{*}=R,R^{\infty},$
for any sequence $\{\rho_{k}\}\subset\mathfrak{S}(\mathcal{H}_{AB})$ converging to a state $\rho_{0}\in\widetilde{\mathfrak{S}}_{II}^{1}(\mathcal{H}_{AB})$.
B) Let $\{\rho_{k}\}$ be an arbitrary sequence of states in $\mathfrak{S}(\mathcal{H}_{AB})$ converging to a state $\rho_{0}\in\widetilde{\mathfrak{S}}_{II}^{1}(\mathcal{H}_{AB})$. The limit relations
$\lim_{k\to+\infty}R^{*}(\rho_{k})=R^{*}(\rho_{0})<+\infty,\quad R^{*}=R,R^{\infty},$ (87)
hold provided that one of the following conditions is valid: [^23] The finiteness of $E_{R}(\rho_{0})$ and $\,I(A\!:\!B)_{\rho_{0}}$ follows from the condition $\rho_{0}\in\widetilde{\mathfrak{S}}_{II}^{1}(\mathcal{H}_{AB})$.
  1. $\rm a)$
    there exists a sequence $\{\omega_{k}\}\subset\mathfrak{T}_{R}$ converging to an operator $\omega_{0}\in\mathfrak{T}_{R}$ such that
  2. $\rm b)$
    there exists
    $\lim_{k\to+\infty}E_{R}(\rho_{k})=E_{R}(\rho_{0})<+\infty;$ (89)
  3. $\rm c)$
    there exists
    $\lim_{k\to+\infty}I(A\!:\!B)_{\rho_{k}}=I(A\!:\!B)_{\rho_{0}}<+\infty.$ (90)
Remark 15. A criterion and different sufficient conditions for the validity of $\,\rm b)$ are described in [7]. Easy verifiable sufficient conditions for the validity of $\,\rm c)$ are presented in [37, Section 4.2].
[…]
B) We prove first that (87) holds provided that condition (88) is valid. By part A it suffices to show that (88) implies that
$\limsup_{k\to+\infty}R^{*}(\rho_{k})\leq R^{*}(\rho_{0}),\quad R^{*}=R,R^{\infty}.$ (91)
The validity of (91) the case $R^{*}=R$ can be shown modifying the arguments from the proof of Theorem 1 in [7].
For given $\lambda\in(0,1)$ and $k\geq 0$ consider the quantity
[…]
The convexity of $\mathfrak{T}_{R}$ implies that $Y_{\lambda,k}(\rho_{k})\geq R(\rho_{k})$, because $(1-\lambda)\sigma+\lambda\omega_{k}\in\mathfrak{T}_{R}$ for all $k$. On the other hand, by the joint convexity of the relative entropy we have
$\begin{array}[]{c}\displaystyle Y_{\lambda,k}\leq(1-\lambda)\inf_{\sigma\in\mathfrak{T}_{R}}(D(\rho_{k}\hskip 1.0pt\|\hskip 1.0pt\sigma)+1-\mathrm{Tr}\sigma)+\lambda(D(\rho_{k}\hskip 1.0pt\|\,\omega_{k})+1-\mathrm{Tr}\omega_{k})\\ \\ =(1-\lambda)R(\rho_{k})+\lambda(D(\rho_{k}\hskip 1.0pt\|\,\omega_{k})+1-\mathrm{Tr}\omega_{k}).\end{array}$
Since we may assume that $D(\rho_{k}\|\hskip 1.0pt\omega_{k})$ is finite for all $k$, we have $R(\rho_{k})=\inf_{\lambda\in(0,1)}Y_{\lambda,k}$ for all $k\geq 0$. Thus, to prove (91) with $R^{*}=R$ it suffices [^24] We use the well-known fact that the infimum of a family of upper semicontinuous functions is an upper semicontinuous function. A more detailed justification for this step can be found in the proof of Theorem 1 in [7]. to show that
$\limsup_{k\to+\infty}Y_{\lambda,k}\leq Y_{\lambda,0}\quad\forall\lambda\in(0,1).$
Since $\mathrm{Tr}\omega_{k}$ tends to $\mathrm{Tr}\omega_{0}$ as $k\to+\infty$, this can be done by proving that
$\lim_{k\to+\infty}D(\rho_{k}\hskip 1.0pt\|\hskip 1.0pt(1-\lambda)\sigma+\lambda\omega_{k})=D(\rho_{0}\hskip 1.0pt\|\hskip 1.0pt(1-\lambda)\sigma+\lambda\omega_{0})<+\infty$
for any operator $\sigma$ in $\mathfrak{T}_{R}$ and any given $\lambda\in(0,1)$. This limit relation follows from (88) by Proposition 2 in [60], because $\lambda\omega_{k}\leq(1-\lambda)\sigma+\lambda\omega_{k}$ for all $k$.
Now it is not hard to show that (88) implies the validity of (91) with $R^{*}=R^{\infty}$.
Let $p\in\mathbb{N}$ be arbitrary. It follows from (88) that
$\lim_{k\to+\infty}D(\rho^{\otimes p}_{k}\|\hskip 1.0pt\omega^{\otimes p}_{k})=D(\rho^{\otimes p}_{0}\|\hskip 1.0pt\omega^{\otimes p}_{0})<+\infty.$
By the above part of the proof this limit relation implies that
$\limsup_{k\to+\infty}R(\rho^{\otimes p}_{k})\leq R(\rho^{\otimes p}_{0})<+\infty.$
Hence, (91) with $R^{*}=R^{\infty}$ holds, because $R^{\infty}(\rho_{k})=\inf_{p\in\mathbb{N}}p^{-1}R(\rho_{k}^{\otimes p})$ for all $k\geq 0$.
Assume that condition (89) holds. Then for an arbitrary subsequence $\{\rho_{k_{t}}\}$ of $\{\rho_{k}\}$ Theorem 1B in [7] guarantees the existence of a sequence of $\{\omega_{t}\}$ of separable states in $\mathfrak{S}(\mathcal{H}_{AB})$ converging to a separable state $\omega_{0}\in\mathfrak{S}(\mathcal{H}_{AB})$ such that
[…]
By the above part of the proof this implies that
$\lim_{t\to+\infty}R^{*}(\rho_{k_{t}})=R^{*}(\rho_{0}),\quad R^{*}=R,R^{\infty},$
because $\{\omega_{t}\}_{t\geq 0}\subset\mathfrak{T}_{R}$. Since the subsequence $\{\rho_{k_{t}}\}$ is arbitrary, we obtain (87).
If condition (90) holds then condition (89) is valid by Proposition 2 in [7]. So, the above part of the proof implies the validity of (87). $\Box$
Remark 16. The proof of Proposition 4.2.2B shows that the limit relations in (91) hold for an arbitrary sequence $\{\rho_{k}\}$ of states in $\mathfrak{S}(\mathcal{H}_{AB})$ converging to any state $\rho_{0}$ in $\mathfrak{S}(\mathcal{H}_{AB})$ provided that one of the conditions $\rm a)$,$\,\rm b)$ and $\rm c)$ is valid.
Remark 17. The properties of the functions $R$ and $R^{\infty}$ established of Proposition 4.2.2 coincide with the properties of the function $E_{R}^{\infty}$ given by Proposition 4.1.3 in Section 4.1.3 in the bipartite case $n=2$.
The convexity of the function $R$ follows from its definition and the joint convexity of the relative entropy, the convexity of the function $R^{\infty}$ follows from the convexity and subadditivity of $R$ by Proposition 13 in [49]. In contrast to $E_{R,\pi}$ and $E_{R,\pi}^{\infty}$, the measurability of $R$ and $R^{\infty}$ is an open question (since $R$ is defined via the infimum over a non-countable set of states). So, to prove the inequalities
$R^{*}(\bar{\rho}(\mu))\leq\int_{\mathfrak{S}(\mathcal{H}_{AB})}\,R^{*}(\varrho)\mu(d\varrho),\quad R^{*}=R,R^{\infty},$ (92)
for a Borel probability measure $\mu$ on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ (where $\bar{\rho}(\mu)$ is the barycenter of $\mu$ defined in (4)) we have first to verify the existence of the integrals in their right hand sides.
Since $R,R^{\infty}\in II_{2}^{1}$, the $II_{m}^{n}$-part of Corollary 3.2 in Section 3 implies the following
Proposition 10.  The functions $R$ and $R^{\infty}$ are $\mu$-integrable w.r.t. a Borel probability measure $\mu$ on $\mathfrak{S}(\mathcal{H}_{AB})$ provided that its barycenter $\bar{\rho}(\mu)$ (defined in (4)) belongs to the set $\,\widetilde{\mathfrak{S}}^{1}_{II}(\mathcal{H}_{AB})$ defined in (86). In this case both inequalities in (92) hold.
4.2.3 Other results about the functions $R$ and $R^{\infty}$
Finite-dimensional approximation of $R$ and $R^{\infty}$. Proposition 4.2.2 in the previous subsection and Lemma 4.1.4 in Section 4.1.4 allows us to prove the following
Proposition 11.  Let $\rho$ be a state from the set $\,\widetilde{\mathfrak{S}}^{1}_{II}(\mathcal{H}_{AB})$ defined in (86). Let $\rho_{k}=Q_{k}\rho\hskip 1.0ptQ_{k}[\mathrm{Tr}Q_{k}\rho\hskip 1.0pt]^{-1}$, $Q_{k}=P_{k}^{A}\otimes P_{k}^{B}$, where $\{P_{k}^{X}\}\subset\mathfrak{B}(\mathcal{H}_{X})$ is an arbitrary sequences of projectors strongly converging to the unit operator $I_{X}$, $X=A,B$. Then
$R^{*}(\rho)=\lim_{k\to+\infty}R^{*}(\rho_{k}),\quad R^{*}=R,R^{\infty}.$
Energy-constrained definition of $R$. Assume that $H_{AB}$ is the Hamiltonian of an infinite-dimensional bipartite system $AB$ and
$\mathfrak{T}^{H_{\!AB}}_{R}=\left\{\sigma\in\mathfrak{T}(\mathcal{H}_{AB})\,|\,\sigma\geq 0,\|\sigma^{\Gamma}\|_{1}\leq 1,\mathrm{Tr}H_{AB}\sigma<+\infty\right\}$
is the subset of $\mathfrak{T}_{R}$ consisting of operators with finite ”mean energy”. Then we may defined the energy-constrained version of the Rains bounds as follows
$R^{H_{\!AB}}(\rho)=\inf_{\sigma\in\mathfrak{T}^{H_{\!AB}}_{R}}(D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma)+1-\mathrm{Tr}\sigma)$
Using the arguments from the proof of Proposition 4.1.5 in Section 4.1.5 one can prove the following
Proposition 12.  Let $H_{A}$ and $H_{B}$ be positive operators on the spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ treated as Hamiltonians of the subsystems $A$ and $B$. If at least one of the operators $H_{A}$ and $H_{B}$ satisfies condition (17) and $H_{AB}=H_{A}\otimes I_{B}+I_{A}\otimes H_{B}$ then
$R^{H_{\!AB}}(\rho)=R(\rho)$
for any state $\rho$ in $\mathfrak{S}(\mathcal{H}_{AB})$ such that $\,\mathrm{Tr}H_{AB}\rho=\mathrm{Tr}H_{A}\rho_{A}+\mathrm{Tr}H_{B}\rho_{B}<+\infty$.
4.3 Conditional entanglement of mutual information
4.3.1 Definitions and basic properties
The conditional entanglement of mutual information is an additive entanglement measure in a multipartite quantum system $A_{1}...A_{n}$ with an operational meaning in terms of the process of partial state merging [61]. For a state $\rho$ of a finite-dimensional system $A_{1}...A_{n}$ it is defined as
$E_{I}(\rho)=\frac{1}{2}\displaystyle\inf_{\hat{\rho}\in\mathfrak{M}_{1}(\rho)}\left[I(A_{1}A^{\prime}_{1}\!:...:\!A_{n}A^{\prime}_{n})_{\hat{\rho}}-I(A^{\prime}_{1}\!:...:\!A^{\prime}_{n})_{\hat{\rho}}\right],$
where $\mathfrak{M}_{1}(\rho)$ is the set of all extensions $\hat{\rho}\in\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}A^{\prime}_{1}..A^{\prime}_{n}})$ of the state $\rho$ [61, 62]. To generalize this definition to an arbitrary state $\rho$ of an infinite-dimensional multipartite system $A_{1}...A_{n}$ note that the quantity $I(A_{1}A^{\prime}_{1}\!:...:\!A_{n}A^{\prime}_{n})_{\varrho}-I(A^{\prime}_{1}\!:...:\!A^{\prime}_{n})_{\varrho}$ well defined for any state $\varrho$ with finite $I(A^{\prime}_{1}\!:...:\!A^{\prime}_{n})_{\varrho}$ coincides at any such state with the nonnegative lower semicontinuous function
$\Delta(\varrho)=I(A_{1}\!:\!A^{\prime}_{2}...A^{\prime}_{n}|A^{\prime}_{1})_{\varrho}+\sum_{k=2}^{n}I(A_{k}\!:\!A_{1}...A_{k-1}A^{\prime}_{1}...A^{\prime}_{k-1}A^{\prime}_{k+1}...A^{\prime}_{n}|A^{\prime}_{k})_{\varrho}$ (93)
well defined for any state $\varrho$ state of the infinite-dimensional system $A_{1}...A_{n}A^{\prime}_{1}...A^{\prime}_{n}$, where all the summands are the extended tripartite QCMI introduced in [16, Theorem 2]. Expression (93) is obtained in [14] in the finite-dimensional settings and generalized to the infinite-dimensional case in [16, Proposition 8].
It is mentioned in [15, Section IV-C] that the function $f=\Delta$ belongs to the class $L_{2n}^{n}(2,n+1)$ (in the notation used in [15]). It follows that the function
$E_{I}(\rho)=\frac{1}{2}\inf_{\hat{\rho}\in\mathfrak{M}_{1}(\rho)}\Delta(\hat{\rho}),$ (94)
on the set of all states of the infinite-dimensional system $A_{1}...A_{n}$ belongs to the class $N^{n}_{n,1}(1,(n+1)/2)\subset II_{n}^{n}$. Faithful $II_{n}^{n}$-type continuity bounds for this function are obtained in [15, Proposition 22]. If fact, a stronger claim is valid.
Lemma 6. The function $E_{I}$ defined in (94) has a faithful $I_{n}^{n}$-type continuity bound, i.e. it belongs to the class $I_{n}^{n}$.
Proof. Let $H_{A_{1}}$,….,$H_{A_{n}}$ be arbitrary positive operators on $\mathcal{H}_{A_{1}}$,….,$\mathcal{H}_{A_{n}}$ satisfying condition (15). By using the continuity bound for the QCMI presented in [26, Lemma 25] it is easy to show the existence of a faithful $I_{n}^{n}$-type continuity bound for the function $\Delta$ defined in (93), i.e. such a function $\,\widehat{B}_{\Delta}(\varepsilon,E|H_{A_{1}},...,H_{A_{n}})\,$ that
$\lim_{\varepsilon\to 0^{+}}\widehat{B}_{\Delta}(\varepsilon,E|H_{A_{1}},...,H_{A_{n}})=0\quad\forall E>0$
and $\,|\Delta(\rho)-\Delta(\sigma)|\leq\widehat{B}_{\Delta}(\varepsilon,E|H_{A_{1}},...,H_{A_{n}})\,$ for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ satisfying the conditions
$\sum_{i=1}^{n}\mathrm{Tr}H_{A_{i}}\rho_{A_{i}},\sum_{i=1}^{n}\mathrm{Tr}H_{A_{i}}\sigma_{A_{i}}\leq nE\quad\textrm{and}\quad\frac{1}{2}\|\rho-\sigma\|_{1}\leq\varepsilon.$ (95)
By using the isometrical equivalence of all purifications of a given state one can show (see Section IV in [33]) that
$E_{I}(\rho)=\frac{1}{2}\inf_{\Lambda}\hskip 1.0pt\Delta(\mathrm{Id}_{A_{1}...A_{n}}\otimes\Lambda(\bar{\rho})),$ (96)
where $\bar{\rho}$ is any pure state in $\,\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}R})$ such that $\mathrm{Tr}_{R}\hskip 1.0pt\bar{\rho}=\rho$ and the infimum is over all channels $\Lambda:\mathfrak{T}(\mathcal{H}_{R})\rightarrow\mathfrak{T}(\mathcal{H}_{A^{\prime}_{1}..A^{\prime}_{n}})$.
Let $\rho$ and $\sigma$ be states in $\,\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ satisfying the conditions in (95). Then there exist pure states $\bar{\rho}$ and $\bar{\sigma}$ in $\,\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}R})$ such that $\,\frac{1}{2}\|\hskip 1.0pt\bar{\rho}-\bar{\sigma}\|_{1}\leq\sqrt{2\varepsilon}$, $\mathrm{Tr}_{R}\hskip 1.0pt\bar{\rho}=\rho$ and $\mathrm{Tr}_{R}\hskip 1.0pt\bar{\sigma}=\sigma$ [2, 9]. Since for any quantum channel $\Lambda:\mathfrak{T}(\mathcal{H}_{R})\rightarrow\mathfrak{T}(\mathcal{H}_{A^{\prime}_{1}..A^{\prime}_{n}})$ we have $\frac{1}{2}\|\mathrm{Id}_{A_{1}...A_{n}}\otimes\Lambda(\bar{\rho})-\mathrm{Id}_{A_{1}...A_{n}}\otimes\Lambda(\bar{\sigma)}\|_{1}\leq\sqrt{2\varepsilon}$ by monotonicity of the trace norm under action of a channel, by applying the $I_{n}^{n}$-type continuity bound $\widehat{B}_{\Delta}(\varepsilon,E|H_{A_{1}},...,H_{A_{n}})$ mentioned before we obtain
$|\Delta(\mathrm{Id}_{A_{1}...A_{n}}\otimes\Lambda(\bar{\rho}))-\Delta(\mathrm{Id}_{A_{1}...A_{n}}\otimes\Lambda(\bar{\sigma}))|\leq\widehat{B}_{\Delta}(\sqrt{2\varepsilon},E|H_{A_{1}},...,H_{A_{n}})$
for any channel $\Lambda:\mathfrak{T}(\mathcal{H}_{R})\rightarrow\mathfrak{T}(\mathcal{H}_{A^{\prime}_{1}..A^{\prime}_{n}})$. So, it follows from (96) that
$|E_{I}(\rho)-E_{I}(\sigma)|\leq\frac{1}{2}\widehat{B}_{\Delta}(\sqrt{2\varepsilon},E|H_{A_{1}},...,H_{A_{n}}).$
The r.h.s. of this inequality tends to zero as $\varepsilon$ tends to zero. So, it is a faithful $I_{n}^{n}$-type continuity bound for the function $E_{I}$. $\Box$
4.3.2 Continuity and convexity properties
The results of Section 3 allows us to prove the following
Proposition 13.  Let $A_{1},..,A_{n}$ be arbitrary quantum systems and
$\mathfrak{S}_{I}^{n}(\mathcal{H}_{A_{1}...A_{n}})=\left\{\hskip 1.0pt\rho\in\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})\,|\,S(\rho_{A_{s}})<+\infty\textrm{ for all }\hskip 1.0pts=\overline{1,n}\hskip 1.0pt\right\}$ (97)
be the subset of $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of states with finite marginal entropies.
A) The function $E_{I}$ is finite and lower semicontinuous on the set $\mathfrak{S}_{I}^{n}(\mathcal{H}_{A_{1}...A_{n}})$. Moreover,
$\liminf_{k\to+\infty}E_{I}(\rho_{k})\geq E_{I}(\rho_{0})$
for any sequence $\{\rho_{k}\}\subset\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ converging to a state $\rho_{0}\in\mathfrak{S}_{I}^{n}(\mathcal{H}_{A_{1}...A_{n}})$.
B) If $\,\{\rho_{k}\}$ is a sequence of states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ converging to a state $\rho_{0}\in\mathfrak{S}_{I}^{n}(\mathcal{H}_{A_{1}...A_{n}})$ such that
$\lim_{k\to+\infty}I(A_{1}\!:...:\!A_{n})_{\rho_{k}}=I(A_{1}\!:...:\!A_{n})_{\rho_{0}}$
then
$\lim_{k\to+\infty}E_{I}(\rho_{k})=E_{I}(\rho_{0})<+\infty.$
Proof. A) By Lemma 4.3.1 the function $E_{I}$ belongs to the class $I_{n}^{n}$. So, this claim follows directly from claim $\rm(i)$ of the $I_{n}^{m}$-part of Corollary 3.2 in Section 3.2, since the function $E_{I}$ does not increase under action of local channels.
B) The convex function $E_{I}$ satisfies the local monotonicity condition (36) and does not exceed the QMI $I(A_{1}\!:...:\!A_{n})$ (because it does not not exceed the c-squashed entanglement $E_{\rm c\textup{-}sq}$ [62]). Thus, to prove this claim it suffices, by Corollary 2 in Section 3.2, to prove that
$\limsup_{k\to+\infty}E_{I}(\rho_{k})\leq E_{I}(\rho_{0})$ (98)
assuming that
$\lim_{k\to+\infty}S([\rho_{k}]_{A_{i}})=S([\rho_{0}]_{A_{i}})<+\infty\quad\textrm{ for all }\;i=\overline{1,n}.$ (99)
For a given $\epsilon>0$ let $\hat{\rho}_{0}$ be a state in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}A^{\prime}_{1}...A^{\prime}_{n}})$ such that $\frac{1}{2}\Delta(\hat{\rho}_{0})\leq E_{I}(\rho_{0})+\epsilon$ (the finiteness of $E_{I}(\rho_{0})$ follows from claim A).
Let $\bar{\rho}_{0}$ be a pure state in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}A^{\prime}_{1}...A^{\prime}_{n}R})$, where $R$ is an infinite-dimensional quantum system, such that $\mathrm{Tr}_{R}\bar{\rho}_{0}=\hat{\rho}_{0}$. The well known results of the purification theory (cf. [2, 9]) imply the existence of a sequence $\{\bar{\rho}_{k}\}$ of pure states in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}A^{\prime}_{1}...A^{\prime}_{n}R})$ converging to the state $\bar{\rho}_{0}$ such that $\mathrm{Tr}_{A^{\prime}_{1}...A^{\prime}_{n}R}\bar{\rho}_{k}=\rho_{k}$ for all $k$. Since the sequence of states $\hat{\rho}_{k}\doteq\mathrm{Tr}_{R}\bar{\rho}_{k}$ in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}A^{\prime}_{1}...A^{\prime}_{n}})$ converges to the state $\hat{\rho}_{0}$, the assumed validity of (99) for all $i=\overline{1,n}$ implies that
$\lim_{k\to+\infty}\Delta(\hat{\rho}_{k})=\Delta(\hat{\rho}_{0})$
by Proposition 8 in [16]. So, it follows from the definition of $E_{I}$ that
$\limsup_{k\to+\infty}E_{I}(\rho_{k})\leq\frac{1}{2}\lim_{k\to+\infty}\Delta(\hat{\rho}_{k})=\frac{1}{2}\Delta(\hat{\rho}_{0})\leq E_{I}(\rho_{0})+\epsilon$
by the construction of the state $\hat{\rho}_{0}$. This implies (98) because $\epsilon$ is arbitrary. $\Box$
The convexity of the function $E_{I}$ in the infinite-dimensional composite system $A_{1}...A_{n}$ is proved by a simple upgrading the finite-dimensional arguments from [61]. Namely, one should establish the following identity
$\Delta\left(\sum_{i\in I}p_{i}\varrho_{i}\otimes|\psi^{1}_{i}\rangle\langle\psi^{1}_{i}|\otimes...\otimes|\psi^{n}_{i}\rangle\langle\psi^{n}_{i}|\right)=\sum_{i\in I}p_{i}\Delta(\varrho_{i})$ (100)
valid for any (finite or countable) set $\{\varrho_{i}\}_{i\in I}$ of states in $\mathfrak{S}(\mathcal{H}_{A_{1}A^{\prime}_{1}...A_{n}A_{n}^{\prime}})$ and any probability distribution $\{p_{i}\}_{i\in I}$, where $\Delta$ in the r.h.s. is the function defined in (93), while $\Delta$ in the l.h.s is the function defined in (93) with $A^{\prime}_{1}$,…,$A_{n}^{\prime}$ replaced, respectively, by $A^{\prime}_{1}A^{\prime\prime}_{1}$,…,$A_{n}^{\prime}A^{\prime\prime}_{1}$, $\{\psi^{1}_{i}\}_{i\in I}$,…$\{\psi^{n}_{i}\}_{i\in I}$ are orthonormal bases in the auxiliary spaces $\mathcal{H}_{A^{\prime\prime}_{1}}$,…,$\mathcal{H}_{A^{\prime\prime}_{n}}$ (their dimension coincides with the cardinality of $I$). If all the states $\rho_{i}$ have finite marginal entropies and the Shannon entropy of the probability distribution $\{p_{i}\}_{i\in I}$ is finite then (100) follows from the coincidence of the r.h.s. of (93) with $\,I(A_{1}A^{\prime}_{1}\!:...:\!A_{n}A^{\prime}_{n})_{\varrho}-I(A^{\prime}_{1}\!:...:\!A^{\prime}_{n})_{\varrho}$. The validity of (100) in the general case is proved via approximation using Proposition 8B in [16].
Using (100) the convexity of $E_{I}$ is proved by simple arguments based on the definition of $E_{I}$ (because the argument of $\Delta$ in the l.h.s. of (100) is an extension of $\sum_{i\in I}p_{i}\rho_{i}$ provided that $\varrho_{i}$ is an extension of $\rho_{i}$). Moreover, since the identity (100) holds also for a countable set of indexes $I$, by this way one can prove the $\sigma$-convexity of $E_{I}$, i.e. the inequality
$E_{I}\left(\sum_{i=1}^{+\infty}p_{i}\rho_{i}\right)\leq\sum_{i=1}^{+\infty}p_{i}E_{I}(\rho_{i})$
for any countable set $\{\rho_{i}\}_{i=1}^{+\infty}$ of states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ and any probability distribution $\{p_{i}\}_{i=1}^{+\infty}$.
It is not clear how to prove the ”continuous” Jensen inequality for the function $E_{I}$ in general. Moreover, it is not clear how to prove that this function is measurable w.r.t. the Borel $\sigma$-algebra in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$. However, claim $\rm(v)$ of the $I_{n}^{m}$-part of Corollary 3.2 in Section 3.2 and the above Lemma 4.3.1 allow us to prove the following
Proposition 14.  Let $A_{1},..,A_{n}$ be arbitrary quantum systems. The function $E_{I}$ defined in (94) is $\mu$-integrable w.r.t. a Borel probability measure $\mu$ on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ provided that its barycenter $\bar{\rho}(\mu)$ (defined in (4)) belongs to the set $\mathfrak{S}^{n}_{I}(\mathcal{H}_{A_{1}...A_{n}})$ defined in (97). In this case
[…]